# Learning Ising Models under Hard Constraints using One Sample

### v1

@@ 3.2 Analysis of First Moment
The lemma below establishes the upper bound on $\phi_{1}(\beta;\bm{\sigma})$. To demonstrate an upper bound on $\phi_{1}(\beta;\bm{\sigma})$ in probability, we use the technique of exchangeable pairs Cha (07) to construct a bound on its variance. With the variance controlled, we invoke Markov’s inequality to conclude $\phi_{1}(\beta;\bm{\sigma})$ that concentrates around its mean.
@@ 3.3.2 Bounding the Magnetizations
$\operatorname{\mathbb{E}}_{\beta^{*}}[m_{i}(\bm{\sigma})^{2}|\bm{\sigma}_{-j}]\geq\frac{\min_{\kappa\in\{-1,1\}}\mathbb{Pr}_{\beta^{*}}[\sigma_{j}=\kappa]}{2}$
@@
In this paper, we present a affirmative answer to the challenge of single sample learning in the truncated Ising model, at all temperatures $\beta\in\mathcal{O}(1)$, giving a sufficient condition for the truncation set $S$, to ensure consistent inference, and extending the existing framework beyond boolean product distributions. Towards this goal, we craft concentration inequalities for the first and second derivatives of the log-pseudolikelihood via arguments concerning the local connectivity of the truncation set.
The present work opens the door to important future questions : (i) Given the above stipulations on the bounded degree $k$-SAT for inference, how does this relate to the discrete distributions constrained to the solutions of a random $k$-SAT? (ii) Is there a way to simultaneously remove the $o(n^{1/6})$ assumption on the maximum degree of the graph while improving the rate of consistency? (iii) Do the current techniques generalize beyond the truncated Ising distribution to include other models, such as truncated Boolean spin glasses?
@@ Appendix A Related Work and Additional Background
The Ising model originated as a mathematical model of ferromagnetism on subgraphs of the lattice $\mathbb{Z}^{d}$, capturing local interactions in physical systems. Ising solved the one-dimensional case in his thesis [34], while Onsager later resolved the two-dimensional case [42], revealing a continuous phase transition between ferromagnetic and paramagnetic states. Beyond low dimensions, the Ising model also serves as a foundational example of spin glasses, aiding both condensed matter physics and probability theory in understanding complex magnetic materials and high-dimensional correlated loss landscapes [46, 47].
Our inquiry into the Ising model will be statistical in nature, concerning the consistent estimation of the inverse temperature parameter under the presence of truncation using a single sample. Despite the seeming simplicity of this task, the presence of phase transitions yields it to be theoretically impossible in certain regimes; our results stand in light of these challenges. One of the primary difficulties in our task is our graph $G$, by extension interaction tensor $A$, is arbitrary (although of a somewhat bounded degree), and thus our results hold in a regime that is neither fully locally connected or mean-field. The first work to prove such a result was [9], who via use of the technique of exchangeable pairs, a variation on Stein’s inequality to prove variance bounds, was able to demonstrate the consistency of the maximum pseudolikelihood estimate derived from a single sample(an objective that will expounded on in the sequel) given the log partition function $F_{G,\beta,n}=\log(Z_{G,\beta,n})$ is diverges with $n$ in the large data limit. As an example, when this seemingly innoucous assumption is not upheld, under the mean field Curie-Weiss model, i.e.
@@
consistent estimation is impossible, as simple calculus yields that $\lim_{n\to\infty}F_{G,\beta,n}=\mathcal{O}(1)$. Moreover, if $\beta$ diverges to infinity with $n$, the psuedolikelihood objective ceases to be strongly concave, collapsing the Fisher information, and rendering estimation impossible. Beyond the estimation of the inverse temperature, [40] was able to extend the regime of Chatterjee [9], demonstrating results for the joint estimation of the inverse temperature $\beta$ and the external field $h$. Viewing the task of estimating the inverse temperature as structure estimation over a parametric class of interaction matrices, i.e parameterized by $\beta$, [11] generalized this setting to provide learning guarantees for large classes of parametric spaces, relying on a clever use of conditioning to use Dobrushin’s condition at all constant temperatures.
@@
A parallel line of work has also investigated the feasibility of learning Markov random fields under hard-constrained distributions with a single sample. This line of work commenced with [6] studying the fugacity parameter of the hard core model, i.e. a probability distribution over independent sets over a graph $G=(V,E)$ represented by binary vectors $\bm{\sigma}\in\{0,1\}^{n}$, where $\sigma_{i}=1$ indicates the node is included in the independent set
@@
Following up on this (and closer to our setting), [27] considered the feasibility of learning boolean product distributions over truncated portions of the boolean hypercube, making use of the tilted-$k$-SAT model over $S\subset\{0,1\}^{n}$, where $S$ is defined to be the set of solutions of a fixed bounded degree $k$-CNF formula $\Phi$.
@@ Appendix B Algorithm For Maximizing the Pseudo-Likelihood
In this section, we present a polynomial-time algorithm for optimizing the pseudo-likelihood objective using projected gradient descent. To guarantee convergence to the optimum, we rely on the following lemma from [1].
Lemma B.1 ([1] Theorem 3.10).
@@
To begin, we demonstrate our upper bound in probability over the first derivative of the log-pseudolikelihood $\phi(\beta;\bm{\sigma})$, showing this concentration inequality via the technique of exchangeable pairs introduced by [9]. Define the anti-symmetric function, $F:S\times S\to\mathbb{R}$,
@@ E.1.1 The Algorithm
Lemma E.2 ([22] Section 23.9).
@@
Lastly, to guarantee there are a sufficient number of flippable variables in the independent set itself, we use the following result from [27] to find a lower bound on the number of variables within the independent set that are flippable under a product distribution.
Lemma E.5 (Lemma 15 & 16 [27]).
@@
The second half of Lemma 3.5 then follows from this corresponding results in [27]. To relate these results to our setting, observe that under our product distribution, the probability that any variable is set to one is at most $e^{-B}/(1+e^{-B})$, and the true marginals may in fact be more balanced. Consequently, the conditions of their result are satisfied in our regime.
@@
We aim to prove this concentration inequality via the technique of exchangeable pairs introduced by [9]. Consider, again, the anti-symmetric function, $F:S\times S\to\mathbb{R}$,

### v2

@@ 3.2 Analysis of First Moment
The lemma below establishes the upper bound on $\phi_{1}(\beta;\bm{\sigma})$. To demonstrate an upper bound on $\phi_{1}(\beta;\bm{\sigma})$ in probability, we use the technique of exchangeable pairs chatterjee2007estimation to construct a bound on its variance. With the variance controlled, we invoke Markov’s inequality to conclude $\phi_{1}(\beta;\bm{\sigma})$ that concentrates around its mean.
@@ 3.3.2 Bounding the Magnetizations
$\operatorname{\mathbb{E}}_{\beta^{*}}[m_{i}(\bm{\sigma})^{2}|\bm{\sigma}_{-j}]\geq\frac{\exp(-B)}{\Delta^{2}}\mathbb{Pr}_{\beta^{*}}[e_{j}(\bm{\sigma})=1]$
@@
In this paper, we present an affirmative answer to the challenge of single sample learning in the truncated Ising model, at all temperatures $\beta\in\mathcal{O}(1)$, giving a sufficient condition for the truncation set $S$ to ensure consistent inference, and extending the existing framework beyond boolean product distributions. Towards this goal, we craft concentration inequalities for the first and second derivatives of the log-pseudolikelihood via local connectivity arguments over the domain.
The present work opens the door to important future questions : (i) Given the above framework, does analyzing measures with random Hamiltonians, like those of the Sherrington-Kirkpatrick model, alleviate the dependence on $\Delta$? (ii) Do logistic regression techniques used to estimate untruncated graphical models apply to the constrained setting? (iii) Is there a way to simultaneously remove the $o(n^{1/6})$ assumption on the maximum degree of the graph while improving the rate of consistency?
@@ Appendix A Related Work and Additional Background
The Ising model originated as a mathematical model of ferromagnetism on subgraphs of the lattice $\mathbb{Z}^{d}$, capturing local interactions in physical systems. Ising solved the one-dimensional case in his thesis [ising1925beitrag], while Onsager later resolved the two-dimensional case [onsager1944crystal], revealing a continuous phase transition between ferromagnetic and paramagnetic states. Beyond low dimensions, the Ising model also serves as a foundational example of spin glasses, aiding both condensed matter physics and probability theory in understanding complex magnetic materials and high-dimensional correlated loss landscapes [34, 35].
Our inquiry into the Ising model will be statistical in nature, concerning the consistent estimation of the inverse temperature parameter under the presence of truncation using a single sample. Despite the seeming simplicity of this task, the presence of phase transitions yields it to be theoretically impossible in certain regimes; our results stand in light of these challenges. One of the primary difficulties in our task is our graph $G$, by extension interaction tensor $A$, is arbitrary (although of a somewhat bounded degree), and thus our results hold in a regime that is neither fully locally connected or mean-field. The first work to prove such a result was [chatterjee2007estimation], who via use of the technique of exchangeable pairs, a variation on Stein’s inequality to prove variance bounds, was able to demonstrate the consistency of the maximum pseudolikelihood estimate derived from a single sample(an objective that will expounded on in the sequel) given the log partition function $F_{G,\beta,n}=\log(Z_{G,\beta,n})$ is diverges with $n$ in the large data limit. As an example, when this seemingly innoucous assumption is not upheld, under the mean field Curie-Weiss model, i.e.
@@
consistent estimation is impossible, as simple calculus yields that $\lim_{n\to\infty}F_{G,\beta,n}=\mathcal{O}(1)$. Moreover, if $\beta$ diverges to infinity with $n$, the psuedolikelihood objective ceases to be strongly concave, collapsing the Fisher information, and rendering estimation impossible. Beyond the estimation of the inverse temperature, [30] was able to extend the regime of Chatterjee [chatterjee2007estimation], demonstrating results for the joint estimation of the inverse temperature $\beta$ and the external field $h$. Viewing the task of estimating the inverse temperature as structure estimation over a parametric class of interaction matrices, i.e parameterized by $\beta$, [6] generalized this setting to provide learning guarantees for large classes of parametric spaces, relying on a clever use of conditioning to use Dobrushin’s condition at all constant temperatures.
@@
A parallel line of work has also investigated the feasibility of learning Markov random fields under hard-constrained distributions with a single sample. This line of work commenced with [bhattacharya2021parameterestimationundirectedgraphical] studying the fugacity parameter of the hard core model, i.e. a probability distribution over independent sets over a graph $G=(V,E)$ represented by binary vectors $\bm{\sigma}\in\{0,1\}^{n}$, where $\sigma_{i}=1$ indicates the node is included in the independent set
@@
Following up on this (and closer to our setting), [galanis2024learning] considered the feasibility of learning boolean product distributions over truncated portions of the boolean hypercube, making use of the tilted-$k$-SAT model over $S\subset\{0,1\}^{n}$, where $S$ is defined to be the set of solutions of a fixed bounded degree $k$-CNF formula $\Phi$.
@@ Appendix B Algorithm For Maximizing the Pseudo-Likelihood
In this section, we present a polynomial-time algorithm for optimizing the pseudo-likelihood objective using projected gradient descent. To guarantee convergence to the optimum, we rely on the following lemma from [bubeck2015convex].
Lemma B.1 ([bubeck2015convex] Theorem 3.10).
@@
To begin, we demonstrate our upper bound in probability over the first derivative of the log-pseudolikelihood $\phi(\beta;\bm{\sigma})$, showing this concentration inequality via the technique of exchangeable pairs introduced by [chatterjee2007estimation]. Define the anti-symmetric function, $F:S\times S\to\mathbb{R}$,
@@ E.1.1 The Algorithm
Lemma E.2 ([frieze2015introduction] Section 23.9).
@@
Lastly, to guarantee there are a sufficient number of flippable variables in the independent set itself, we use the following result from [galanis2024learning] to find a lower bound on the number of variables within the independent set that are flippable under a product distribution.
Lemma E.5 (Lemma 15 & 16 [galanis2024learning]).
@@
The second half of Lemma 3.5 then follows from this corresponding results in [galanis2024learning]. To relate these results to our setting, observe that under our product distribution, the probability that any variable is set to one is at most $e^{-B}/(1+e^{-B})$, and the true marginals may in fact be more balanced. Consequently, the conditions of their result are satisfied in our regime.
@@
We aim to prove this concentration inequality via the technique of exchangeable pairs introduced by [chatterjee2007estimation]. Consider, again, the anti-symmetric function, $F:S\times S\to\mathbb{R}$,